# A Compartmental Model for Epidemiology with Human Behavior and Stochastic Effects

### v2

@@ 3. An SIR model with noncompliant behavior and stochastic perturbation
In essence, we would like to replace $\beta$ with $\beta+\sigma_{\beta}\dot{W}$ where $\sigma_{\beta}\geq 0$ is the level of uncertainty in $\beta$. Similarly, we would like to replace $\mu$ with $\mu+\sigma_{\mu}\dot{W}$ for some level of uncertainty $\sigma_{\mu}\geq 0$. However, for technical reasons involving the proof of nonnegativity and global existence of solutions, it is desirable to ensure that the uncertainty in any given equation is proportional to the population which the equation describes. Thus we opt for the following dynamics:
(3.1) $\begin{split}dS&=((1-\xi)b-\beta(1-\alpha)SI^{(M)}-\mu SN^{*}+\nu S^{*}-\delta S)dt+(-\sigma_{\beta}(1-\alpha)^{2}SI-\sigma_{\mu}SS^{*})dW\\ dI&=(\beta(1-\alpha)SI^{(M)}-\gamma I-\mu IN^{*}+\nu I^{*}-\delta I)dt+(\sigma_{\beta}(1-\alpha)^{2}SI-\sigma_{\mu}II^{*})dW,\\ dR&=(\gamma I-\delta R-\mu RN^{*}+\nu R^{*})dt-\sigma_{\mu}RR^{*}dW,\\ dS^{*}&=(\xi b-\beta S^{*}I^{(M)}+\mu SN^{*}-\nu S^{*}-\delta S^{*})dt+(-\sigma_{\beta}S^{*}I^{*}+\sigma_{\mu}SS^{*})dW,\\ dI^{*}&=(\beta S^{*}I^{(M)}-\gamma I^{*}+\mu IN^{*}-\nu I^{*}-\delta I^{*})dt+(\sigma_{\beta}S^{*}I^{*}+\sigma_{\mu}II^{*})dW,\\ dR^{*}&=(\gamma I^{*}-\delta R^{*}+\mu RN^{*}-\nu R^{*})dt+\sigma_{\mu}RR^{*}dW.\end{split}$
From a modeling perspective, one may prefer to replace each of $SI,S^{*}I^{*}$ in the stochastic terms with $SI^{(M)},S^{*}I^{(M)}$ (and similarly for stochastic terms involving $\sigma_{\mu}$). We make these particular modeling decisions so that that stochastic term in each equation is proportional to the population itself, which is necessary in (3.6) in the proof of theorem 3.1 below.
@@
Since all terms in (3.1) are locally Lipschitz with respect to $(S,I,R,S^{*},I^{*},R^{*})$, given positive initial data, a local solution of (3.1) exists on a time interval $[0,\tau_{E})$. We’d like to prove that this solution remains bounded (hence exists globally) and nonnegative. To do so, starting with $k_{0}\in{\mathbb{N}}$ such that
@@
where $F_{A}$ represents the underlying dynamics and $\Sigma_{A}$ represents the stochastic drift. By Itô’s formula [28, Ch. 2], we have
@@
Operating on the time interval $[0,\tau_{k})$, we now simplify and bound (3.3) term-by-term, remembering that populations are nonnegative on this time interval. In what follows $C$ is a positive constant that changes from line-to-line and depends on the ambient parameters $b,\delta,\alpha,\beta,\gamma,\mu,\nu.$ Looking at (3.1), we have
@@
[…]
Since $f$ is bounded for times in $[0,\tau_{k})$, in expectation the stochastic integral disappears and we have
@@ 4. Asymptotic behavior of the stochastic system near disease free states
In this section, we perform some asymptotic analysis of (3.1) with respect to certain disease free states. There are two disease free equilibria for the stochastic system in different parameter regimes. The first is $X_{1}=(\frac{b}{\delta},0,0,0,0,0)$ in the case that $\xi=0$, which is addressed in the deterministic regime in theorem 2.5(i). The second is $X_{2}=(0,0,0,\tfrac{b}{\delta},0,0)$ in the case that $\xi=1$ and $\nu=0$. As discussed in remark 2.6, this latter scenario can be thought of as the “worst case scenario” DFE and is addressed as part of theorem 2.5(iii).
Besides these, we would like to say something about the behavior of solutions to (3.1) near the disease free states $X_{s,s^{*}}=(s,0,0,s^{*},0,0)$ where $s,s^{*}$ are the non-zero elements of $x_{3}$ as defined in (2.21). These are equilibrium points for the deterministic system (2.2) and are addressed in theorem 2.5(ii) when $\xi=0$ and $\tfrac{b}{\delta}>\tfrac{\delta+\nu}{\mu}$ and in theorem 2.5(iii) when $\xi\in(0,1]$. These are not equilibrium points for the stochastic system (3.1) due to the stochastic drift terms. However, we would like to quantify how far the stochastic drift can push the dynamics away from these states if the stability conditions of 2.5(ii) or 2.5(iii) are met.
Our first two stability results follow from lemma 4.1 which is a special case of Theorem 4.4 in [22, Ch. 4]. Before introducing the lemma, we establish some preliminary notation. Consider the stochastic ODE in $d$-dimensions
@@
By Itô’s formula [28, Ch. 2], assuming $X$ satisfies (4.1), we have
@@
[22, Ch. 4, Thm. 4.4] Assume that there is a non-negative function $V\in C^{2}(R^{d})$ and positive constants $A,B,C$, such that
@@
For the following theorem, we assume that $\xi=0$ and define the basic reproduction number corresponding to the fully compliant disease free equilibrium $X_{1}=(\frac{b}{\delta},0,0,0,0,0)$ by
@@
~$\tilde{\ell}_{3}=(v+v^{*})\Big{(}(1-\alpha)\Big{(}u+\tfrac{b}{\delta}\Big{)}+u^{*}\Big{)}((1-\alpha)v+v^{*}).$
@@
(4.11) ~$\begin{split}\tilde{\ell}_{3}&\leq\tfrac{b}{\delta}(1-\alpha)(v+v^{*})((1-\alpha)v+v^{*})+\alpha u^{*}(v+v^{*})((1-\alpha)v+v^{*})\\ &\leq\tfrac{b}{\delta}(1-\alpha)((1-\alpha)v^{2}+(2-\alpha)vv^{*}+(v^{*})^{2})+\tfrac{\alpha b}{\delta}((1-\alpha)u^{*}v+u^{*}v^{*}).\end{split}$
@@
(4.12) ~$\tilde{\ell}_{3}\leq\tfrac{b}{\delta}((1-\alpha)^{2}+\varepsilon)v^{2}+\tfrac{C}{\varepsilon}(u^{*})^{2}+\tfrac{C}{\varepsilon}(v^{*})^{2}$
@@
Next we turn our attention to worst case scenario of $\xi=1$ and $\nu=0.$ In this case, the DFE is $X_{2}=(0,0,0,\tfrac{b}{\delta},0,0)$ and we define the reproductive ratio by
@@
Suppose $\xi=1$ and $\nu=0.$ If $\mathscr{R}^{\sigma}_{0}(0,\tfrac{b}{\delta})<1,$ then there are positive constants $C_{1},C_{2}$ such that for any solution $(S,I,R,S^{*},I^{*},R^{*})$ of (3.1), we have
@@
We split theorem 4.4 into two statements to reflect what might actually be important to a policy-maker. Here $X_{2}$ could be unstable in two different ways: (1) infections could increase and persist or (2) the stochasticity in transmission of noncompliance (represented by $\sigma_{\mu}$) could be large enough that on average, enough noncompliant individuals are randomly acquiring compliance so that compliance persists. For the policy-maker, scenario (2) could be seen as a benefit, so there is no reason to actually hope that $X_{2}$ is stable with respect to this sort of perturbation. However, one would want to guarantee that scenario (1) does not occur. The first statement, regarding decay of second moments of $I$ and $I^{*}$, gives conditions under which the disease is expected to die out. The second statement regarding mean-square stability is included for mathematical completeness, though its unclear if this is even desirable in a real-world scenario. The condition $\frac{\sigma_{\mu}^{2}}{2}\Big{(}\tfrac{b}{\delta}\Big{)}^{2}<\delta$ in the second statement guarantees that the compliant population does not grow (on average). As seen in (4.27) below, this could be relaxed to $\frac{\sigma_{\mu}^{2}}{2}\Big{(}\tfrac{b}{\delta}\Big{)}^{2}<\delta+\eta\mu$ where $\eta$ is any lower bound on the total noncompliant population. Regardless, setting $\sigma_{\beta}=\sigma_{\mu}=0$, the latter conditions reduces to $\delta>0,$ and $\mathscr{R}^{\sigma}_{0}(0,\tfrac{b}{\delta})$ reduces to $\mathscr{R}_{0}(0,\tfrac{b}{\delta})$ as defined in (2.16), and thus theorem 4.4 reduces to the sufficient stability condition of theorem 2.5(iii) with $\xi=1$ and $\nu=0$.
@@
[…]
@@
If in addition we assume that $\Big{(}\tfrac{b}{\delta}\Big{)}^{2}\sigma_{\mu}^{2}<\delta$, then we can prove $X_{2}=(0,0,0,\tfrac{b}{\delta},0,0)$ is exponentially mean-square stable by finding a suitable Lyapunov function in the same manner as in the proof of theorem 4.2. In this case, letting $u=S,v=I,w=R,u^{*}=S^{*}-\tfrac{b}{\delta},v^{*}=I^{*},w^{*}=R^{*}$, we define
@@
[22, Ch. 1, Thm. 3.4] Suppose that $M=M(t)$ is a scalar, real-valued, continuous martingale vanishing at $t=0$. Then
[…]
@@
We note that $V_{2}\geq 0$ for $n^{*}\geq-s^{*}$ (indeed, if we treat $V_{2}$ as a function of $n^{*}$, we find $V_{2}^{\prime\prime}\geq 0$ and $V_{2}^{\prime}=V_{2}=0$ at $n^{*}=0$, implying that $V_{2}$ takes a minimum value of 00 at $n^{*}=0$). Also
@@
Here we have discarded $-4\delta vv^{*}$ which is the only cross term known to be nonpositive. We use the inequality $-2(\gamma+2\delta)uv\leq\tfrac{\delta}{2}u^{2}+\tfrac{2(\gamma+2\delta)^{2}}{\delta}v^{2}$ and similarly for the other cross terms involving the constant $(\gamma+2\delta).$ This results in
@@
The last thing to do is to deal with the $v$ and $v^{*}$ equations. To do this, we define $V_{6}=v^{2}+(v^{*})^{2}=\|(v,v^{*})\|^{2}_{2}.$ Using the linearization (2.13), we see that
@@
where $\Sigma_{v}=\sigma_{\beta}(1-\alpha)^{2}(u+s)v-\sigma_{\mu}vv^{*}$, $\Sigma_{v^{*}}=\sigma_{\beta}(u^{*}+s^{*})v^{*}+\sigma_{\mu}vv^{*}.$ In particular, we use $\Sigma_{v}^{2},\Sigma_{v^{*}}^{2}\leq 2\Big{(}\tfrac{b}{\delta}\Big{)}^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})$. Thus we have
(4.43) $\begin{split}dV_{6}&\leq\bigg{(}2\left\langle\begin{bmatrix}v\\ v^{*}\end{bmatrix},(F_{s,s^{*}}-V_{s,s_{*}})\begin{bmatrix}v\\ v^{*}\end{bmatrix}\right\rangle\\ &\hskip 28.45274pt+2\left\langle\begin{bmatrix}v\\ v^{*}\end{bmatrix},\begin{bmatrix}\beta(1-\alpha)u((1-\alpha)v+v^{*})-\mu vn^{*}\\ \beta u^{*}((1-\alpha)v+v^{*})+\mu vn^{*}\end{bmatrix}\right\rangle\\ &\hskip 28.45274pt+2\Big{(}\tfrac{b}{\delta}\Big{)}^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg{)}dt+fdW\end{split}$
Since we are assuming $\mathscr{R}_{0}(s,s^{*})<1$, we know that all eigenvalues of $F_{s,s^{*}}-V_{s,s^{*}}$ have negative real part [33], and thus
$\left\langle\begin{bmatrix}v\\ v^{*}\end{bmatrix},(F_{s,s^{*}}-V_{s,s_{*}})\begin{bmatrix}v\\ v^{*}\end{bmatrix}\right\rangle\leq-2\varepsilon(v^{2}+(v^{*})^{2})$
for some positive $\varepsilon$. For the other inner product term, we simply add positive terms to arrive at bounds
$u((1-\alpha)v+v^{*})\leq n((1-\alpha)v+v^{*}),\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,u^{*}((1-\alpha)v+v^{*})\leq n^{*}((1-\alpha)v+v^{*}).$
Then using $v,v^{*}\leq\tfrac{b}{\delta}$, each term in the inner product is bounded by a constant multiple of $vn,vn^{*},v^{*}n$ or $v^{*}n^{*}$. Passing all the constants on to $n,n^{*}$, one particular bound we can achieve is
$2\left\langle\begin{bmatrix}v\\ v^{*}\end{bmatrix},\begin{bmatrix}\beta(1-\alpha)u((1-\alpha)v+v^{*})-\mu vn^{*}\\ \beta u^{*}((1-\alpha)v+v^{*})+\mu vn^{*}\end{bmatrix}\right\rangle\leq\varepsilon(v^{2}+(v^{*})^{2})+\tfrac{6\beta+3\mu}{2\varepsilon}(n^{2}+(n^{*})^{2})$
so that (4.43) becomes
(4.44) $dV_{6}\leq\bigg{(}-\varepsilon(v^{2}+(v^{*})^{2})+\tfrac{6\beta+3\mu}{2\varepsilon}(n^{2}+(n^{*})^{2})+2\Big{(}\tfrac{b}{\delta}\Big{)}^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg{)}dt+fdW.$
@@
(4.45) $\begin{split}dV&\leq\bigg{(}-\delta(u^{2}+(u^{*})^{2}+w^{2}+(w^{*})^{2})+\Big{(}\tfrac{6\beta+3\mu}{2\varepsilon}-2c_{1}\delta\Big{)}(n^{2}+(n^{*})^{2})\\ &\hskip 28.45274pt+\Big{(}2\delta-\tfrac{4\gamma^{2}+4(\gamma+2\delta)^{2}+2\beta^{2}(2\delta/\nu)}{\delta}+c_{2}\varepsilon\Big{)}(v^{2}+(v^{*})^{2})\\ &\hskip 28.45274pt+C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg{)}dt+fdW\end{split}$
Choosing $c_{1}=\tfrac{6\beta+3\mu}{4\varepsilon\delta}$ and $c_{2}=\tfrac{4\gamma^{2}+4(\gamma+2\delta)^{2}+2\beta^{2}(2\delta/\nu)}{\varepsilon\delta}$, we arrive at
(4.46) $dV\leq(-\delta(u^{2}+v^{2}+w^{2}+(u^{*})^{2}+(v^{*})^{2}+(w^{*})^{2}+C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2}))dt+fdW.$
@@
[…]

### v4

@@ 3. An SIR model with noncompliant behavior and stochastic perturbation
In essence, we would like to replace $\beta$ with $\beta+\sigma_{\beta}\dot{W}$ where $\sigma_{\beta}\geq 0$ is the level of uncertainty in $\beta$. Similarly, we would like to replace $\mu$ with $\mu+\sigma_{\mu}\dot{W}$ for some level of uncertainty $\sigma_{\mu}\geq 0$. However, for technical reasons involving the proof of nonnegativity and global existence of solutions, it is desirable to ensure that the uncertainty in any given equation is proportional to the population which the equation describes. Thus we consider the following dynamics:
(3.1) $\begin{split}dS&=((1-\xi)b-\beta(1-\alpha)SI^{(M)}-\mu SN^{*}+\nu S^{*}-\delta S)dt+(-\sigma_{\beta}(1-\alpha)^{2}SI-\sigma_{\mu}SS^{*})dW,\\ dI&=(\beta(1-\alpha)SI^{(M)}-\gamma I-\mu IN^{*}+\nu I^{*}-\delta I)dt+(\sigma_{\beta}(1-\alpha)^{2}SI-\sigma_{\mu}II^{*})dW,\\ dR&=(\gamma I-\mu RN^{*}+\nu R^{*}-\delta R)dt-\sigma_{\mu}RR^{*}dW,\\ dS^{*}&=(\xi b-\beta S^{*}I^{(M)}+\mu SN^{*}-\nu S^{*}-\delta S^{*})dt+(-\sigma_{\beta}S^{*}I^{*}+\sigma_{\mu}SS^{*})dW,\\ dI^{*}&=(\beta S^{*}I^{(M)}-\gamma I^{*}+\mu IN^{*}-\nu I^{*}-\delta I^{*})dt+(\sigma_{\beta}S^{*}I^{*}+\sigma_{\mu}II^{*})dW,\\ dR^{*}&=(\gamma I^{*}+\mu RN^{*}-\nu R^{*}-\delta R^{*})dt+\sigma_{\mu}RR^{*}dW.\end{split}$
From a modeling perspective, one may prefer to replace each of $(1-\alpha)^{2}SI,S^{*}I^{*}$ in the stochastic terms with $(1-\alpha)SI^{(M)},S^{*}I^{(M)}$ (and similarly for stochastic terms involving $\sigma_{\mu}$). We make these particular modeling decisions so that that stochastic term in each equation is proportional to the population itself, which is necessary in (3.6) in the proof of theorem 3.1 below. We note that, besides the proof of theorem 3.1, this choice has no effect, so that theorems of section 4 still hold for the full system where $(1-\alpha)^{2}SI,S^{*}I^{*}$ in the stochastic terms are replaced with $(1-\alpha)SI^{(M)},S^{*}I^{(M)}$, with the caveat that one must assume global existence of a solution which is eventually nonnegative. This is because in the proof of theorems 4.2 and 4.4, the only bound that we use for the stochastic terms is the total population bound $\tfrac{b}{\delta}$ which works just as well for the full stochastic terms as for the reduced ones. Further, with or without this modification, system (3.1) captures the same behavior in the asymptotic regimes studied in section 4 since those regimes alternately involve the whole population becoming compliant so that $(1-\alpha)SI^{(M)}\to(1-\alpha)^{2}SI$, the whole population becoming noncompliant so that $S^{*}I^{(M)}\to S^{*}I^{*}$, and similar observations can be made regarding the stochastic terms involving spread of noncompliance.
@@
Since all terms in (3.1) are locally Lipschitz with respect to $(S,I,R,S^{*},I^{*},R^{*})$, given positive initial data, a local solution of (3.1) exists on a time interval $0,\tau_{E})$. We’d like to prove that this solution remains bounded (hence exists globally) and nonnegative. To do so, starting with $k_{0}\in{\mathbb{N}}$ such that
@@
where $F_{A}$ represents the underlying dynamics and $\Sigma_{A}$ represents the stochastic drift. By Itô’s formula [47, Ch. 2], we have
@@
Operating on the time interval $0,\tau_{k})$, we now simplify and bound (3.3) term-by-term, remembering that populations are nonnegative on this time interval. In what follows $C$ is a positive constant that changes from line-to-line and depends on the ambient parameters $b,\delta,\alpha,\beta,\gamma,\mu,\nu.$ Looking at (3.1), we have
@@
(3.8) $\int^{t}_{0}dV\leq\int^{t}_{0}Mdt+\int^{t}_{0}f(S,I,R,S^{*},I^{*},R^{*})dW.$
Since $f$ is bounded for times in $0,\tau_{k})$, in expectation the stochastic integral disappears and we have
@@ 4. Asymptotic behavior of the stochastic system near disease free states
In this section, we perform some asymptotic analysis of (3.1) with respect to certain disease free states. In particular, we prove stability results for (3.1) analogous to 2.5(i) and the “worst case scenario” DFE addressed as part of theorem 2.5(iii) and discussed in remark 2.6. Besides these, we would like to say something about the behavior of solutions to (3.1) near the disease free states $X_{s,s^{*}}=(s,0,0,s^{*},0,0)$ where $s,s^{*}$ are the non-zero elements of $x_{3}$ as defined in (2.21). These are not equilibria for the stochastic model but we would like to quantify how far the stochastic drift can push the dynamics away from these states if the stability conditions of theorem 2.5(ii) or 2.5(iii) are met.
Our first two stability results follow from lemma 4.1 which is a special case of Theorem 4.4 in [41, Ch. 4]. Before introducing the lemma, we establish some preliminary notation. Consider the stochastic ODE in $d$-dimensions
@@
By Itô’s formula [47, Ch. 2], assuming $X$ satisfies (4.1), we have
@@
[41, Ch. 4, Thm. 4.4] Assume that there is a non-negative function $V\in C^{2}(R^{d})$ and positive constants $A,B,C$, such that
@@
For the following theorem, we assume that $\xi=0$ and define a disease extinction threshold corresponding to the fully compliant disease free equilibrium $X_{1}=(\frac{b}{\delta},0,0,0,0,0)$ by
@@
$\tilde{\ell}_{3}=(v+v^{*})\Big((1-\alpha)\Big(u+\tfrac{b}{\delta}\Big)+u^{*}\Big)((1-\alpha)v+v^{*}).$
@@
(4.11) $\begin{split}\tilde{\ell}_{3}&\leq\tfrac{b}{\delta}(1-\alpha)(v+v^{*})((1-\alpha)v+v^{*})+\alpha u^{*}(v+v^{*})((1-\alpha)v+v^{*})\\ &\leq\tfrac{b}{\delta}(1-\alpha)((1-\alpha)v^{2}+(2-\alpha)vv^{*}+(v^{*})^{2})+\tfrac{\alpha b}{\delta}((1-\alpha)u^{*}v+u^{*}v^{*}).\end{split}$
@@
(4.12) $\tilde{\ell}_{3}\leq\tfrac{b}{\delta}((1-\alpha)^{2}+\varepsilon)v^{2}+\tfrac{C}{\varepsilon}(u^{*})^{2}+\tfrac{C}{\varepsilon}(v^{*})^{2}$
@@
Next we turn our attention to worst case scenario of $\xi=1$ and $\nu=0.$ In this case, the DFE is $X_{2}=(0,0,0,\tfrac{b}{\delta},0,0)$ and we define a disease extinction threshold by
@@
Suppose $\xi=1$ and $\nu=0.$ If $\mathscr{R}^{\sigma}_{0}(0,\tfrac{b}{\delta})<1,$ then for any solution $(S,I,R,S^{*},I^{*},R^{*})$ of (3.1), we have
@@
We split theorem 4.4 into two statements to reflect what might actually be important to a policy-maker. Here $X_{2}$ could be unstable in two different ways: (1) infections could increase and persist or (2) the stochasticity in transmission of noncompliance could cause random spikes in compliance to occur. For the policy-maker, scenario (2) could be seen as a benefit, so there is no reason to actually hope that $X_{2}$ is stable with respect to this sort of perturbation. However, one would want to guarantee that scenario (1) does not occur. The first statement, regarding decay of second moments of $I$ and $I^{*}$, gives conditions under which the disease is expected to die out. The second statement regarding mean-square stability is included for mathematical completeness, though its unclear if this is even desirable in a real-world scenario. The condition $\frac{\sigma_{\mu}^{2}}{2}\Big(\tfrac{b}{\delta}\Big)^{2}<\delta$ in the second statement guarantees that the compliant population does not randomly spike. As seen in (4.27) below, this could be relaxed to $\frac{\sigma_{\mu}^{2}}{2}\Big(\tfrac{b}{\delta}\Big)^{2}<\delta+\eta\mu$ where $\eta$ is any lower bound on the total noncompliant population. Regardless, setting $\sigma_{\beta}=\sigma_{\mu}=0$, the latter conditions reduces to $\delta>0,$ and $\mathscr{R}^{\sigma}_{0}(0,\tfrac{b}{\delta})$ reduces to $\mathscr{R}_{0}(0,\tfrac{b}{\delta})$ as defined in (2.16), and thus theorem 4.4 reduces to the sufficient stability condition of theorem 2.5(iii) with $\xi=1$ and $\nu=0$.
@@
with $C$ as defined in (4.22), which is negative due to the assumption that $\mathscr{R}_{0}^{\sigma}(0,\tfrac{b}{\delta})<1$. The function $f$ is bounded, since its arguments remain bounded, and thus $\mathbb{E}\Big(\int^{t}_{0}fdW\Big)=0$, so integrating (4.25), taking the expectation and applying Fubini’s theorem, we arrive at
$\mathbb{E}(V)\leq V_{0}-C\int^{t}_{0}\mathbb{E}(V)dt.$
@@
If in addition we assume that $\frac{\sigma_{\mu}^{2}}{2}\Big(\tfrac{b}{\delta}\Big)^{2}<\delta$, then we can prove $X_{2}=(0,0,0,\tfrac{b}{\delta},0,0)$ is exponentially mean-square stable by finding a suitable Lyapunov function in the same manner as in the proof of theorem 4.2. In this case, letting $u=S,v=I,w=R,u^{*}=S^{*}-\tfrac{b}{\delta},v^{*}=I^{*},w^{*}=R^{*}$, we define
@@
[41, Ch. 1, Thm. 3.4] Suppose that $M=M(t)$ is a scalar, real-valued, continuous martingale vanishing at $t=0$. Then
$\limsup_{t\to\infty}\frac{M(t)^{2}}{t}<\infty\,\,\,\text{a.s.}\,\,\,\,\,\,\implies\,\,\,\,\,\,\limsup_{t\to\infty}\frac{M(t)}{t}=0\,\,\,\text{a.s.}$
@@
We note that $V_{2}\geq 0$ for $n^{*}\geq-s^{*}$ (indeed, if we treat $V_{2}$ as a function of $n^{*}$, we find $V_{2}^{\prime\prime}\geq 0$ and $V_{2}^{\prime}=V_{2}=0$ at $n^{*}=0$, implying that $V_{2}$ takes a minimum value of 0 at $n^{*}=0$). Also
@@
Here we have discarded $4\delta vv^{*}$ which is the only cross term known to be nonpositive. We use the inequality $2(\gamma+2\delta)uv\leq\tfrac{\delta}{2}u^{2}+\tfrac{2(\gamma+2\delta)^{2}}{\delta}v^{2}$ and similarly for the other cross terms involving the constant $(\gamma+2\delta).$ This results in
@@
The last thing to do is to deal with the $v$ and $v^{*}$ equations. Using the linearization (2.13), we see that
@@
where $\Sigma_{v}=\sigma_{\beta}(1-\alpha)^{2}(u+s)v-\sigma_{\mu}vv^{*}$, $\Sigma_{v^{*}}=\sigma_{\beta}(u^{*}+s^{*})v^{*}+\sigma_{\mu}vv^{*}.$ In particular, we have $\Sigma_{v}^{2},\Sigma_{v^{*}}^{2}\leq 2\Big(\tfrac{b}{\delta}\Big)^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})$. Since we are assuming $\mathscr{R}_{0}(s,s^{*})<1$, we know that all eigenvalues of $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=F_{s,s^{*}}-V_{s,s^{*}}$ have negative real part [53]. Borrowing from the theory of Sylvester equations [11, Ch. 5],[14, Ch. 8], this is a necessary and sufficient condition for the existence of a symmetric positive definite matrix $Q$ such that $M^{t}Q+QM=-I$. We define
(4.43) $V_{6}=\left\langle Q\begin{bmatrix}v\\ v^{*}\end{bmatrix},\begin{bmatrix}v\\ v^{*}\end{bmatrix}\right\rangle=\|(v,v^{*})\|^{2}_{Q}.$
A quick computation shows that
(4.44) $\begin{split}dV_{6}&=\bigg(-v^{2}-(v^{*})^{2}+2\left\langle Q\begin{bmatrix}v\\ v^{*}\end{bmatrix},\begin{bmatrix}\beta(1-\alpha)u((1-\alpha)v+v^{*})-\mu vn^{*}\\ \beta u^{*}((1-\alpha)v+v^{*})+\mu vn^{*}\end{bmatrix}\right\rangle\\ &\hskip 113.81102pt+2\left\langle Q\begin{bmatrix}\Sigma_{v}\\ \Sigma_{v^{*}}\end{bmatrix},\begin{bmatrix}\Sigma_{v}\\ \Sigma_{v^{*}}\end{bmatrix}\right\rangle\bigg)dt+fdW\\ &=\bigg(-v^{2}-(v^{*})^{2}+2\left\langle Q\begin{bmatrix}v\\ v^{*}\end{bmatrix},\begin{bmatrix}\beta(1-\alpha)u((1-\alpha)v+v^{*})-\mu vn^{*}\\ \beta u^{*}((1-\alpha)v+v^{*})+\mu vn^{*}\end{bmatrix}\right\rangle\\ &\hskip 113.81102pt+8\|Q\|_{2}\big(\tfrac{b}{\delta})^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg)dt+fdW\end{split}.$
For the remaining inner product term in (4.44), we use $v,v^{*}\leq\tfrac{b}{\delta}$, as well as
(4.45) $(1-\alpha)u((1-\alpha)v+v^{*})\leq n(v+v^{*}),\,\,\,\,\,\,\,\,u^{*}((1-\alpha)v+v^{*})\leq n(v+v^{*}).$
Having done so, all remaining terms are bounded by constant multiples of $vn,v^{*}n,vn^{*}$ or $v^{*}n^{*}$, and we bound each of these using $\beta vn\leq\tfrac{3\beta^{2}n^{2}}{2\varepsilon}+\tfrac{\varepsilon v^{2}}{6}$, and similarly for other terms. All of this leads to the bound
(4.46) $\begin{split}dV_{6}&\leq\bigg(-v^{2}-(v^{*})^{2}+2\|Q\|_{2}\Big(\tfrac{b}{\delta}\Big)\bigg(\tfrac{2\varepsilon}{3}v^{2}+\tfrac{\varepsilon}{3}(v^{*})^{2}+\tfrac{3\beta}{\varepsilon}n^{2}+\tfrac{3(\beta+\mu)}{\varepsilon}(n^{*})^{2}\bigg)\\ &\hskip 113.81102pt+8\|Q\|_{2}\Big(\tfrac{b}{\delta}\Big)^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg)dt+fdW\\ &\leq\bigg(-\tfrac{1}{3}v^{2}-\tfrac{1}{3}(v^{*})^{2}+12\|Q\|^{2}_{2}\Big(\tfrac{b}{\delta}\Big)^{2}(\beta+\mu)(n^{2}+(n^{*})^{2})dt\\ &\hskip 113.81102pt+8\|Q\|_{2}\Big(\tfrac{b}{\delta}\Big)^{4}(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg)dt+fdW,\end{split}$
where we chose $\varepsilon=1/(2\|Q\|_{2}(b/\delta))$ in the last line.
@@
(4.47) $\begin{split}dV&\leq\bigg(-\delta(u^{2}+(u^{*})^{2}+w^{2}+(w^{*})^{2})+\Big(12\|Q\|^{2}_{2}\Big(\tfrac{b}{\delta}\Big)^{2}(\beta+\mu)c_{2}-2c_{1}\delta\Big)(n^{2}+(n^{*})^{2})\\ &\hskip 28.45274pt-\Big(2\delta-\tfrac{4\gamma^{2}+4(\gamma+2\delta)^{2}+2\beta^{2}(2\delta/\nu)}{\delta}+\frac{c_{2}}{3}\Big)(v^{2}+(v^{*})^{2})\\ &\hskip 28.45274pt+C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\bigg)dt+fdW\end{split}$
Choosing $c_{2}=\tfrac{12\gamma^{2}+12(\gamma+2\delta)^{2}+6\beta^{2}(2\delta/\nu)}{\delta}$ and then $c_{1}=\frac{12\|Q\|^{2}_{2}(b/\delta)^{2}(\beta+\mu)c_{2}}{2\delta}$, we arrive at
(4.48) $dV\leq\Big(-\delta(u^{2}+v^{2}+w^{2}+(u^{*})^{2}+(v^{*})^{2}+(w^{*})^{2})+C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\Big)dt+fdW.$
@@
$\frac{V(t)-V(0)}{t\delta}+\frac{1}{t}\int^{t}_{0}\Big(u^{2}+v^{2}+w^{2}+(u^{*})^{2}+(v^{*})^{2}+(w^{*})^{2}\Big)\leq C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})+\frac{1}{t}\int^{t}_{0}fdW.$
Taking the $\limsup$ as $t\to\infty$, the result follows since $V$ is nonnegative and the integral on the right hand side goes to zero almost surely by lemma 4.6. ∎